# Visuomotor Policy Learning via Coarse-to-Fine Autoregressive Prediction

### v2

@@ 1 Introduction
Figure 1: Policy Comparison. The representative performance among Behavior Transformer [52] served as an autoregressive policy, Diffusion Policy [11], and our approach in the state-based Robomimic square task experiment. CARP achieves a superior balance of performance and efficiency.
@@
Policy learning from demonstrations, formulated as the supervised regression task of mapping observations to actions, has proven highly effective across various robotic tasks, even in its simplest form. Replacing the policies with generative models, particularly those stemming from the vision community, has opened new avenues for improving performance, enabling robots to achieve the high precision required for complex tasks.
Existing approaches have explored different generative modeling techniques to address challenges in visuomotor policy learning. Autoregressive Modeling (AM) [52, 70, 12, 30] provides a straightforward and efficient out-of-the-box solution, benefitting from its scalability, flexibility, and mature exploration at lower computation requirements. However, AM’s next-token prediction paradigm often fails to capture long-range dependencies, global structure, and temporal coherence [26], leading to poor performance, which is essential for many robotic tasks. Recently, Diffusion Modeling (DM) [11, 61, 51] has emerged as a promising alternative, bridging the precision gap in AM by modeling the gradient of the action score function to learn multimodal distributions. Nevertheless, DM requires multiple steps of sequential denoising, making it computationally prohibitive for robotic tasks, especially for robotic tasks requiring efficient real-time inference in on-board compute-constrained environments. Additionally, DDPM [19]’s rigid generative process lacks flexibility and adaptability for tasks with long-term dependencies, often leading to cumulative errors and reduced robustness over extended time spans.
Both AM and DM have their respective advantages and limitations, which are often orthogonal and difficult to balance in practical applications. In this work, we aim to resolve this trade-off by introducing a novel generative paradigm for robot visuomotor policy learning that predicts entire action sequences from a coarse-to-fine granularity in a next-scale prediction framework. This approach allows our model to achieve performance levels comparable to DM while achieving AM’s efficiency and flexibility.
Our primary contribution is the introduction of a Coarse-to-Fine AutoRegressive Policy (CARP), a hybrid framework that combines AM’s efficiency with DM’s high performance to meet the demands of real-world robotic manipulation. Specifically, our contributions are as follows:
@@
  • Coarse-to-fine autoregressive prediction: This mechanism refines action sequences in the latent space using Cross-Entropy loss with relaxed Markovian assumptions during iterations, achieving DM-like performance with high efficiency.
@@ 2 Background
We start by introducing the background, focusing on three key areas: problem formulation, conventional autoregressive policies, and diffusion-based policies.
@@ 2.1 Problem Formulation
Problem formulation will consider a task $\mathcal{T}$, where there are $N$ expert demonstrations $\{\tau_{i}\}_{i=1}^{N}$. Each demonstration $\tau_{i}$ is a sequence of state-action pairs. We formulate robot imitation learning as an action sequence prediction problem [46, 11, 61], training a model to minimize the error in future actions conditioned on historical states. Specifically, imitation learning minimize the behavior cloning loss $\mathcal{L}_{bc}$ formulated as
$\mathcal{L}_{bc}=\mathbb{E}_{\boldsymbol{s},\boldsymbol{a}\sim\mathcal{T}}\left[\sum_{t=0}^{T}\mathcal{L}\left(\pi_{\boldsymbol{\theta}}(\boldsymbol{a}_{H}|\boldsymbol{s}_{O}),\boldsymbol{a}_{H}\right)\right],$ (1)
@@ 2.2 Autoregressive Policy
Autoregressive policies leverage the efficiency and flexibility of autoregressive models (GPT-style Decoders). Recent advancements, such as action chunking [52, 70], redefine this paradigm as a multi-token, one-pass prediction method (see Appendix B for further details). We refer to this approach as Autoregressive Policy (AP). In this approach, the next-token posits the probability of observing the current action $\boldsymbol{a}_{t}$ depends solely on its previous states $\boldsymbol{s}_{O}$, which allows for the factorization of the likelihood of sequence with length $H$ as:
$p\left(\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},...,\boldsymbol{a}_{t+H-1}\right)=\prod^{t+H-1}_{k=t}p\left(\boldsymbol{a}_{k}|\boldsymbol{s}_{O}\right).$ (2)
However, it introduces several issues that hinder its performance. The linear, step-by-step unidirectional dependency of action prediction may overlook the global structure [26], making it challenging to capture long-range dependencies and holistic coherence or temporal locality [22] in complex scenes or sequences, which restricts their generalizability in tasks requiring bidirectional reasoning, as shown in Fig. 2(a). For example, it cannot predict the former actions given the near-terminal state.
@@ 2.3 Diffusion Policy
Diffusion-based policies [11] utilize Denoising Diffusion Probabilistic Models [19] to approximate the conditional distribution $p(\boldsymbol{a}_{H}|\boldsymbol{s}_{O})$, through modeling the noise during the denoising process from Gaussian noise to noise-free output as
$\boldsymbol{a}_{H}^{k+1}=\alpha(\boldsymbol{a}_{H}^{k}-\gamma\epsilon_{\theta}(\boldsymbol{s}_{O},\boldsymbol{a}_{H}^{k},k)+\mathcal{N}),$ (3)
where $\epsilon_{\theta}$ is the learnable noise network, $k$ is the current denosing step, $\mathcal{N}$ is the Gaussian noise, $\alpha$ and $\gamma$ are the hyper-parameters.
Diffusion-based policies show impressive performance in robotic manipulation tasks due to their action generation which gradually refines from random samples which we can abstract as a coarse-to-fine process. This kind of process alike human movement or natural thinking flows in line with human intuition. However, they suffer from poor convergence due to their design. To address this, multiple denoising steps are required, constrained by the Markovian assumption, where $\boldsymbol{a}_{H}^{k+1}$ depends only on the previous step $\boldsymbol{a}_{H}^{k}$ (as shown in Fig. 2(b)), which leads to significant runtime inefficiency. Moreover, they lack the ability to leverage generation context effectively, hindering scalability to complex scenes [16].
@@ 3 Method
To address the limitations of existing methods, we propose Coarse-to-Fine AutoRegressive Policy (CARP), a novel visuomotor policy framework that combines the high performance of recent diffusion-based policies with the sample efficiency and flexibility of traditional autoregressive policies.
CARP achieves these advantages through a redesigned autoregressive modeling strategy. Specifically, we shift from conventional next-token prediction to a coarse-to-fine, next-scale prediction approach, using multi-scale action representations. CARP’s training is divided into two stages: multi-scale action tokenization and coarse-to-fine autoregressive prediction, as shown in Fig. 3.
@@ 3.1 Multi-Scale Action Tokenization
Instead of focusing on individual action steps, we extract representations at multiple scales across the entire action sequence. We propose a novel multi-scale action quantization autoencoder that encodes a sequence of actions into $K$ discrete token maps, $\boldsymbol{R}=(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$, which are used for both training and inference. Our approach builds on the VQVAE architecture [60], incorporating a modified multi-scale quantization layer [58] to enable hierarchical encoding.
Encoder and Decoder. As illustrated in Fig. 3(a), the actions are first organized into an action sequence $\boldsymbol{A}^{H\times D}$, where $H$ denotes the prediction horizon and $D$ represents the dimensionality of the action $\boldsymbol{a}$. Given that each dimension in the action space is orthogonal to the others, and that there exists a natural temporal dependency within an action sequence, we employ a 1D temporal convolutional network (1D-CNN) along the time dimension, as used in [22], for both the encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$. Let $\boldsymbol{F}=\mathcal{E}(\boldsymbol{A})\in\mathbb{R}^{L\times C}$, where $L$ denotes the compressed length of the temporal dimension with $L\leq H$, and $C$ represents the dimensionality of the feature map $\boldsymbol{F}$.
Quantization. We introduce a quantizer with a learnable codebook $\mathcal{Z}\in\mathbb{R}^{V\times C}$, containing $V$ code vectors, each of dimension $C$. The quantization process (lines 4-11 in Algorithm 1) generates the action sequence representation iteratively across multiple scales. At scale $k$, it produces action token map $\boldsymbol{r}_{k}$ to represent the sequence, which consists of $l_{k}$ tokens $q$. In our implementation, we set $l_{k}=k$. The function $\operatorname{Lookup}(\mathcal{Z},v)$ retrieves the $v$-th code vector from $\mathcal{Z}$. The quantization function is defined by $\boldsymbol{r}=\mathcal{Q}(\boldsymbol{F})$ as following:
$q=\operatorname*{arg\,min}_{v\in[V]}\operatorname{dist}\left(\operatorname{Lookup}(\mathcal{Z},v),\boldsymbol{f}\right)\in[V],$ (4)
where token $q$ represents the nearest vector in $\mathcal{Z}$ for a given feature vector $\boldsymbol{f}\in\mathbb{R}^{1\times C}$ in the feature map $\boldsymbol{F}$, based on a distance function $\operatorname{dist}(\cdot)$.
@@
Loss. The final approximation $\hat{\boldsymbol{F}}$ of the original feature map $\boldsymbol{F}$ is combined residually with the multi-scale representation derived from the codebook $\mathcal{Z}$, based on each $\boldsymbol{r}$. The reconstructed action sequence is then obtained as $\hat{\boldsymbol{A}}=\mathcal{D}(\hat{\boldsymbol{F}})$. To train the quantized autoencoder, a typical VQVAE loss $\mathcal{L}$ is minimized as following
$\mathcal{L}=\underbrace{\|\boldsymbol{A}-\hat{\boldsymbol{A}}\|_{2}}_{\mathcal{L}_{\text{recon}}}+\underbrace{\|\text{sg}(\boldsymbol{F})-\hat{\boldsymbol{F}}\|_{2}}_{\mathcal{L}_{\text{quant}}}+\underbrace{\|\boldsymbol{F}-\text{sg}(\hat{\boldsymbol{F}})\|_{2}}_{\mathcal{L}_{\text{commit}}},$ (5)
where $\mathcal{L}_{\text{recon}}$ minimizes the difference between the original action sequence $\boldsymbol{A}$ and the reconstruction $\hat{\boldsymbol{A}}$. $\text{sg}(\cdot)$ denotes stop-gradient. $\mathcal{L}_{\text{quant}}$ aligns the quantized feature map $\hat{\boldsymbol{F}}$ with the original $\boldsymbol{F}$, and $\mathcal{L}_{\text{commit}}$ encourages the encoder to commit to codebook entries, preventing codebook collapse. For $\mathcal{L}_{\text{quant}}$ and $\mathcal{L}_{\text{commit}}$, we calculate every residual calculating moment of each scale $\boldsymbol{r}_{k}$ as in Algorithm 1. After training, the autoencoder $\{\mathcal{E},\mathcal{Q},\mathcal{D}\}$ tokenizes actions for subsequent coarse-to-fine autoregressive modeling.
@@
Unlike traditional autoregressive policies that predict each action token independently (see Eq. 2 and Fig. 2(a)), CARP leverages dual capabilities to capture both local temporal dependencies and global structure across the entire action sequence. This approach produces smoother transitions and more stable action sequences. Through multi-scale encoding, CARP overcomes the short-sighted limitations of conventional autoregressive models, yielding more robust, coherent, and precise behaviors over extended time horizons.
@@ 3.2 Coarse-to-Fine Autoregressive Prediction
Prediction. As we receive the multi-scale representation tokens $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$, each $\boldsymbol{r}$ represents the same action sequence in different scales. The autoregressive likelihood can be formulated as
$p(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})=\prod_{k=1}^{K}p(\boldsymbol{r}_{k}\mid\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1};\boldsymbol{s}_{O}),$ (6)
where each autoregressive unit $\boldsymbol{r}_{k}\in[V]^{l_{k}}$ is the token map at scale $k$ containing $l_{k}$ tokens $q$. The predix sequence $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1})$ is served as the condition for $\boldsymbol{r}_{k}$, accompanying with the historical state sequence $\boldsymbol{s}_{O}$. This kind of next-scale prediction methodology is what we define as coarse-to-fine autoregressive prediction.
@@
Loss. During the $k$-th autoregressive step, all distributions over the $l_{k}$ tokens in $\boldsymbol{r}_{k}$ will be generated in parallel, with the coarse-to-fine dependency ensured by a block-wise causal attention mask [58]. To optimize the autoregressive model, we utilize the standard Cross-Entropy loss to capture the difference between the predicted token map $\hat{\boldsymbol{r}}$ and the token map $\boldsymbol{r}$ from the ground truth action sequence:
$\mathcal{L}_{\text{Cross-Entropy}}=\sum_{k=1}^{K}\sum_{i=1}^{l_{k}}\left[-\sum_{v=1}^{V}\boldsymbol{r}_{k}^{i,v}\log\hat{\boldsymbol{r}}_{k}^{i,v}\right],$ (7)
where $l_{k}$ is the length of each scale, $V$ is the action vocabulary size, and $K$ is the number of scales.
Discussion. CARP models the entire trajectory holistically, progressively refining actions from high-level intentions to fine-grained details (see Fig. 2(c)). This approach aligns more closely with natural human behavior, where movement is guided by overarching intentions rather than step-by-step planning.
Instead of the commonly used MSE loss [70, 59, 5], CARP employs Cross-Entropy loss (Eq. 7) in behavior cloning, as MSE tends to enforce a unimodal distribution that can be detrimental to manipulation tasks [52]. CARP’s iterative refinement process resembles the denoising steps in diffusion models to achieve high accuracy.
Furthermore, our approach models actions directly rather than modeling noise, enabling faster convergence in low-dimensional manifolds [36, 64]. By operating in the latent space with action sequence tokens, CARP mitigates trajectory anomalies that can disrupt prediction, unlike methods that work on raw actions [11]. This latent-space representation allows CARP to focus on the essential components of actions, resulting in smoother and more efficient predictions. In contrast, traditional diffusion models rely on Markovian processes, which compress all information into progressively noisier inputs from previous levels, often hindering efficient learning and requiring more inference steps [16] (see Eq. 3 and Fig. 2(b)). CARP relaxes this constraint by allowing each scale $\boldsymbol{r}_{k}$ to depend on all prior scales $\boldsymbol{r}_{1:k-1}$ instead of the only previous scale $\boldsymbol{r}_{k-1}$. This structure enables CARP to generate high-quality trajectories with significantly fewer steps.

### v3

# CARP: Visuomotor Policy Learning via Coarse-to-Fine Autoregressive Prediction

@@ 1 Introduction
Figure 1: Policy Comparison. The representative performance among Behavior Transformer [54] served as an autoregressive policy, Diffusion Policy [12], and our approach in the state-based Robomimic Square task experiment. CARP demonstrates an effective balance between task performance and inference efficiency.
@@
Policy learning from demonstrations, formulated as the supervised regression task of mapping observations to actions, has proven highly effective across a wide range of robotic tasks, even in its simplest form. Replacing the traditional policies with generative models, particularly those stemming from the vision community, has opened new avenues for improving performance, enabling robots to achieve the high precision required for complex tasks in robotic scenarios.
Existing approaches have explored different generative modeling techniques to address challenges in visuomotor policy learning. Autoregressive Modeling (AM) [54, 72, 13, 32] provides a straightforward and efficient out-of-the-box solution, benefitting from its scalability, flexibility, and mature exploration at lower computation requirements. However, AM’s next-token prediction paradigm often fails to capture long-range dependencies, global structure, and temporal coherence [28], leading to poor performance, which is essential for many robotic tasks. Recently, Diffusion Modeling (DM) [12, 63, 53] has emerged as a promising alternative, bridging the precision gap in AM by modeling the gradient of the action score function to learn multimodal distributions. Nevertheless, DM requires multiple steps of sequential denoising, making it computationally prohibitive for robotic tasks, especially for robotic tasks requiring efficient real-time inference in on-board compute-constrained environments. Additionally, DDPM [21]’s rigid generative process lacks flexibility and adaptability for tasks with long-term dependencies, often leading to cumulative errors and reduced robustness over extended time spans, limiting its use in dynamic settings.
Both AM and DM have their respective advantages and limitations, which are often orthogonal and difficult to balance in practical applications. In this work, we aim to resolve this trade-off by introducing a novel generative paradigm for robot visuomotor policy learning that predicts entire action sequences from a coarse-to-fine granularity in a next-scale prediction framework. This approach allows our model to achieve performance levels comparable to DM while retaining AM’s inference efficiency and implementation flexibility.
Our primary contribution is the introduction of a Coarse-to-Fine AutoRegressive Policy (CARP), a hybrid framework that combines AM’s efficiency with DM’s high performance to meet the demands of real-world robotic manipulation. Specifically, our main contributions are as follows:
@@
  • Coarse-to-Fine autoregressive prediction: This mechanism refines action sequences in the latent space using Cross-Entropy loss with relaxed Markovian assumptions during iterations, achieving DM-like performance with high efficiency and comparable multi-modal behavior.
@@ 2 Background
We start by providing the background of our approach, focusing on three key components: problem formulation, conventional autoregressive policies, and diffusion-based policies.
@@ 2.1 Problem Formulation
Problem formulation will consider a task $\mathcal{T}$, where there are $N$ expert demonstrations $\{\tau_{i}\}_{i=1}^{N}$. Each demonstration $\tau_{i}$ is a sequence of state-action pairs. We formulate robot imitation learning as an action sequence prediction problem [48, 12, 63], training a model to minimize the error between the predicted future actions conditioned on historical states and the ground truth actions. Specifically, imitation learning minimize the behavior cloning loss $\mathcal{L}_{bc}$ formulated as
$\mathcal{L}_{bc}=\mathbb{E}_{\boldsymbol{s},\boldsymbol{a}\sim\mathcal{T}}\left[\sum_{t=0}^{T}\mathcal{L}\left(\pi_{\boldsymbol{\theta}}(\hat{\boldsymbol{a}}_{H}|\boldsymbol{s}_{O}),\boldsymbol{a}_{H}\right)\right],\vskip-7.11317pt$ (1)
@@ 2.2 Autoregressive Policy
Autoregressive policies leverage the efficiency and flexibility of autoregressive models (GPT-style Decoders). Recent advancements, such as action chunking [54, 72], redefine this paradigm as a multi-token, one-pass prediction method (see Suppl. C for further details). We refer to this approach as Autoregressive Policy (AP). In this approach, the next-token posits the probability of observing the current action $\boldsymbol{a}_{t}$ depends solely on its previous states $\boldsymbol{s}_{O}$, which allows for the factorization of the likelihood of sequence with length $H$ as
$p\left(\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},...,\boldsymbol{a}_{t+H-1}\right)=\prod^{t+H-1}_{k=t}p\left(\boldsymbol{a}_{k}|\boldsymbol{s}_{O}\right).\vskip-7.11317pt$ (2)
However, it introduces several issues that hinder its performance. The linear, step-by-step unidirectional dependency of action prediction may overlook the global structure [28], making it challenging to capture long-range dependencies and holistic coherence or temporal locality [24] in complex scenes or sequences, which restricts their generalizability in tasks requiring bidirectional reasoning, as shown in Fig. 2(a). For example, it cannot predict the former actions given the near-terminal state without backward reasoning.
@@ 2.3 Diffusion Policy
Diffusion-based policies [12] utilize Denoising Diffusion Probabilistic Models [21] to approximate the conditional distribution $p(\boldsymbol{a}_{H}|\boldsymbol{s}_{O})$ instead of the joint distribution [24] $p(\boldsymbol{a}_{H},\boldsymbol{s}_{O})$, through modeling the noise during the denoising process from Gaussian noise to noise-free output as
$\boldsymbol{a}_{H}^{k+1}=\alpha(\boldsymbol{a}_{H}^{k}-\gamma\epsilon_{\theta}(\boldsymbol{s}_{O},\boldsymbol{a}_{H}^{k},k)+\mathcal{N}),\vskip-7.11317pt$ (3)
where $\epsilon_{\theta}$ is a learnable noise network, $k$ is the current denoising step, $\mathcal{N}$ is Gaussian noise, and $\alpha$, $\gamma$ are hyper-parameters.
Diffusion-based policies show impressive performance in robotic manipulation tasks due to their action generation which gradually refines from random samples which we can abstract as a coarse-to-fine process. This kind of process alike human movement or natural thinking flows in line with human intuition. However, they suffer from poor convergence due to their design. To address this, multiple denoising steps are required, constrained by the Markovian assumption, where $\boldsymbol{a}_{H}^{k+1}$ depends only on the previous step $\boldsymbol{a}_{H}^{k}$ (see Fig. 2(b)) , which leads to significant runtime inefficiency. Moreover, they lack the ability to leverage generation context effectively, hindering scalability to complex scenes [17].
@@ 3 Method
To address the limitations of existing methods, we propose Coarse-to-Fine AutoRegressive Policy (CARP), a novel visuomotor policy framework that combines the high performance of recent diffusion-based policies with the inference efficiency and flexibility of traditional autoregressive policies. CARP achieves these advantages through a redesigned autoregressive modeling strategy. Specifically, we shift from conventional next-token prediction to a coarse-to-fine, next-scale prediction approach, using multi-scale action representations. The training process of CARP consists of two stages: multi-scale action tokenization and coarse-to-fine autoregressive prediction, as detailed in Fig. 3.
@@ 3.1 Multi-Scale Action Tokenization
Instead of focusing on individual action steps, we extract representations at multiple scales across the entire action sequence. We propose a novel multi-scale action quantization autoencoder that encodes a sequence of actions into $K$ discrete token maps, $\boldsymbol{R}=(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$, which are used for training and inference. Our approach builds on the VQVAE architecture [62], incorporating a modified multi-scale quantization layer [60] to enable hierarchical encoding.
Encoder and Decoder. As illustrated in Fig. 3(a), the actions are first organized into an action sequence $\boldsymbol{A}^{H\times D}$, where $H$ denotes the prediction horizon and $D$ represents the dimensionality of the action $\boldsymbol{a}$. Given that each dimension in the action space is orthogonal to the others, and that there exists a natural temporal dependency within an action sequence, we employ a 1D temporal convolutional network (1D-CNN) along the time dimension, similar to the architecture used in [24], for both the encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$. Let $\boldsymbol{F}=\mathcal{E}(\boldsymbol{A})\in\mathbb{R}^{L\times C}$, where $L$ denotes the compressed length of the temporal dimension with $L\leq H$, and $C$ represents the dimensionality of the feature map $\boldsymbol{F}$.
Quantization. We introduce a quantizer with a learnable codebook $\mathcal{Z}\in\mathbb{R}^{V\times C}$, containing $V$ code vectors, each of dimension $C$. The quantization process (lines 4-11 in Algorithm 1) generates the action sequence representation iteratively across multiple scales. At scale $k$, it produces action token map $\boldsymbol{r}_{k}$ to represent the sequence, which consists of $l_{k}$ tokens $q$. In our implementation, we set $l_{k}=k$. The function $\operatorname{Lookup}(\mathcal{Z},v)$ retrieves the $v$-th code vector from $\mathcal{Z}$. The quantization function is defined by $\boldsymbol{r}=\mathcal{Q}(\boldsymbol{F})$ as
$q=\operatorname*{arg\,min}_{v\in[V]}\operatorname{dist}\left(\operatorname{Lookup}(\mathcal{Z},v),\boldsymbol{f}\right)\in[V],\vskip-5.69054pt$ (4)
where token $q$ represents the nearest vector in $\mathcal{Z}$ for a given feature vector $\boldsymbol{f}\in\mathbb{R}^{1\times C}$ in the feature map $\boldsymbol{F}$, according to a distance function $\operatorname{dist}(\cdot)$, such as Euclidean, cosine, etc.
@@
[…]
$\mathcal{L}=\underbrace{\|\boldsymbol{A}-\hat{\boldsymbol{A}}\|_{2}}_{\mathcal{L}_{\text{recon}}}+\underbrace{\|\text{sg}(\boldsymbol{F})-\hat{\boldsymbol{F}}\|_{2}}_{\mathcal{L}_{\text{quant}}}+\underbrace{\|\boldsymbol{F}-\text{sg}(\hat{\boldsymbol{F}})\|_{2}}_{\mathcal{L}_{\text{commit}}},\vskip-7.11317pt$ (5)
where $\mathcal{L}_{\text{recon}}$ minimizes the difference between the original action sequence $\boldsymbol{A}$ and the reconstruction $\hat{\boldsymbol{A}}$. $\text{sg}(\cdot)$ denotes stop-gradient. $\mathcal{L}_{\text{quant}}$ aligns the quantized feature map $\hat{\boldsymbol{F}}$ with the original $\boldsymbol{F}$, and $\mathcal{L}_{\text{commit}}$ encourages the encoder to commit to codebook entries, preventing codebook collapse. For $\mathcal{L}_{\text{quant}}$ and $\mathcal{L}_{\text{commit}}$, we calculate every residual calculating moment of each scale $\boldsymbol{r}_{k}$ as in Algorithm 1. After the training process, the autoencoder $\{\mathcal{E},\mathcal{Q},\mathcal{D}\}$ tokenizes actions for subsequent coarse-to-fine autoregressive modeling.
@@
Unlike traditional autoregressive policies that predict each action token independently (see Eq. 2 and Fig. 2(a)), CARP leverages dual capabilities to capture both local temporal dependencies and global features across the entire action sequence. This approach produces smoother transitions and more stable action sequences. Through multi-scale encoding, CARP overcomes the short-sighted limitations of conventional autoregressive models, yielding more robust, coherent, and precise behaviors over extended time horizons.
@@ 3.2 Coarse-to-Fine Autoregressive Prediction
Prediction. Given the multi-scale representation token maps $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})$ produced by the trained autoencoder, where each $\boldsymbol{r}_{k}$ encodes the same action sequence $\boldsymbol{A}$ at a different scale, the autoregressive likelihood is formulated as
$p(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{K})=\prod_{k=1}^{K}p(\boldsymbol{r}_{k}\mid\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1};\boldsymbol{s}_{O}),\vskip-2.84526pt$ (6)
where each autoregressive unit $\boldsymbol{r}_{k}\in[V]^{l_{k}}$ is the token map at scale $k$ containing $l_{k}$ tokens $q$. The predix sequence $(\boldsymbol{r}_{1},\boldsymbol{r}_{2},\dots,\boldsymbol{r}_{k-1})$ is served as the condition for $\boldsymbol{r}_{k}$, accompanying with the historical state sequence $\boldsymbol{s}_{O}$. This kind of next-scale prediction methodology is what we define as coarse-to-fine autoregressive prediction in CARP.
@@
Loss. During the $k$-th autoregressive step, all distributions over the $l_{k}$ tokens in $\boldsymbol{r}_{k}$ will be generated in parallel, with the coarse-to-fine dependency ensured by a block-wise causal attention mask [60]. To optimize the autoregressive model, we utilize the standard Cross-Entropy loss to capture the difference between the predicted token map $\hat{\boldsymbol{r}}$ and the token map $\boldsymbol{r}$ from the ground truth action sequence as
$\mathcal{L}_{\text{Cross-Entropy}}=\sum_{k=1}^{K}\sum_{i=1}^{l_{k}}\left[-\sum_{v=1}^{V}\boldsymbol{r}_{k}^{i,v}\log\hat{\boldsymbol{r}}_{k}^{i,v}\right],\vskip-7.11317pt$ (7)
where $l_{k}$ is the length of each scale, $V$ is the size of the action dictionary $\mathcal{Z}$, and $K$ is the number of scales.
Discussion. CARP models the entire trajectory holistically, progressively refining actions from high-level intentions to fine-grained details (see Fig. 2(c)). This hierarchical paradigm more closely resembles natural human behavior, where movements are guided by overarching goals and gradually refined, rather than being planned step by step.
Instead of the commonly used MSE loss [72, 61, 6], CARP employs Cross-Entropy loss (Eq. 7) to preserve multimodality naturally, as MSE tends to enforce a unimodal distribution that can be detrimental to robotic tasks [54]. CARP’s iterative refinement process resembles the denoising steps in diffusion models to achieve high accuracy.
Furthermore, our approach models actions directly rather than modeling noise, facilitating faster prediction convergence in low-dimensional manifolds [38, 66]. By operating in the latent space with action sequence tokens, CARP mitigates trajectory anomalies that can disrupt prediction, unlike methods that work on raw actions [12]. This latent-space representation allows CARP to focus on the essential components of actions, resulting in smoother and more efficient predictions. In contrast, traditional diffusion models rely on Markovian processes, which compress all information into progressively noisier inputs from previous levels, often hindering efficient learning and requiring more inference steps [17] (see Eq. 3 and Fig. 2(b)). CARP relaxes this constraint by allowing each scale $\boldsymbol{r}_{k}$ to depend on all prior scales $\boldsymbol{r}_{1:k-1}$ instead of the only previous scale $\boldsymbol{r}_{k-1}$. This structure enables CARP to generate high-quality trajectories with significantly fewer steps, demonstrating superior efficiency.